OpenStudy (anonymous):

i got a tricky one for you guys. bobs friends decided to equally share $270 cost of a cookware set for a wedding gift.however before the wedding 3 more of bobs friends joined the others in giving the couple the cookware. this lowered the contibutuin of each by $15. how many friends ORIGINALLY agreeed to share the cost of it.

OpenStudy (accessdenied):

So, I started with two definitions. \( n := \text{the number of original friends} \), \( C := \text{original amount paid by each friend.} \) We can then create two equations relating \(n\) and \(C\). \( \Large \frac{270}{n} = C \) > Each friend is paying an equal amount, so it is evenly distributed to each friend. \( \Large \frac{270}{n+3} = C - 15 \) > When three more friends join in, the amount paid by each goes down by 15. Then, I substituted the first equation into the second to create a proportion. \( \Large \frac{270}{n+3} = \frac{270}{n} - 15 \) \( \Large \frac{270}{n+3} = \frac{270}{n} - \frac{15n}{n} \) \( \Large \frac{270}{n+3} = \frac{270 - 15n}{n} \) I multiply by the lowest common denominator between the values, n(n+3) on both sides (or, essentially cross multiplication) \( \large 270n = (270 - 15n)(n + 3) \) \( \large 270n = 270n + 810 - 15n^2 - 45n \) \( \large 0 = -15n^2 - 45n + 810 \) We can factor out a -15 and divide it off. \( \large -15(n^2 + 3n - 45) = 0 \)\ \( \large n^2 + 3n - 45 = 0 \) Then factor and solve for \(n > 0\) and \(n\) are not restricted in the original equations: \( \large (n + 9)(n - 6) = 0 \) \( \large n + 9 = 0 \text{ and } n - 6 = 0 \) \( \large \cancel{n = -9} \text{ and } \boxed{n = 6} \) Thus, the original number of friends was \( 6 \text{ friends} \).

OpenStudy (accessdenied):

Of course, as I check, it does appear to work out as well. Letting \(n=6\), \[ \frac{270}{n} = \frac{270}{6} = 45 \] \[ \frac{270}{n+3} = \frac{270}{9} = 30 \] which is 15 less than the original.

OpenStudy (anonymous):

\[\frac{270}{n}-15\text{ = }\frac{270}{n+3}\]n = 6 or -9, n=6